# Spectrum and location of ongoing extreme particle acceleration in Cassiopeia A

### v1

@@ Abstract
Young supernova remnants (SNRs) are believed to be the origin of energetic cosmic rays (CRs) below the “knee” of their spectrum at $\sim 3$ petaelectronvolt (PeV, $10^{15}$ eV). Nevertheless, the precise location, duration, and operation of CR acceleration in young SNRs are open questions. Here, we report on multi-epoch X-ray observations of Cassiopeia A (Cas A), a 350-year-old SNR, in the 15-50 keV band that probes the most energetic CR electrons. The observed X-ray flux decrease $(15\pm 1\%)$, contrary to the expected $>$90% decrease based on previous radio, X-ray, and gamma-ray observations, provides unambiguous evidence for CR electron acceleration operating in Cas A. A temporal model for the radio and X-ray data accounting for electron cooling and continuous injection finds that the freshly injected electron spectrum is significantly harder (exponential cutoff power law index $q=2.15$), and its cutoff energy is much higher ($E_{cut}=36$ TeV) than the relic electron spectrum ($q=2.44\pm 0.03$, $E_{cut}=4\pm 1$ TeV). Both electron spectra are naturally explained by the recently developed modified nonlinear diffusive shock acceleration (mNLDSA) mechanism. The CR protons producing the observed gamma rays are likely accelerated at the same location by the same mechanism as those for the injected electron. The Cas A observations and spectral modeling represent the first time radio, X-ray, gamma ray and CR spectra have been self-consistently tied to a specific acceleration mechanism -mNLDSA– in a young SNR.
@@ 2.1 Observations and data reduction
Table 1: List of observations used in this analysis
@@
40001019002 2012-08-18 84 291
40021001002 2012-08-27 76 170
40021001005 2012-10-07 33 184
40021002002 2012-11-23 338 271
40021002006 2013-03-02 249 135
40021002008 2013-03-05 249 189
40021003003 2013-05-28 151 197
40021011002 2013-10-30 7 235
40021012002 2013-11-27 335 205
40021015002 2013-12-21 312 75
40021015003 2013-12-23 312 137
Total exposure (archival data) 2,089
40801003002 2023-03-24 229 92
40801013002 2023-04-04 217 95
@@ 2.3 Spectral analysis
We analyzed the data in the 15-50 keV band. This choice of energy range ensures negligible contribution of thermal X-rays ($kT$ $<4$ keV (Hwang & Laming, 2012; Bleeker et al., 2001; Maeda et al., 2009b; Favata et al., 1997b; Allen et al., 1997b)) and source emission above the background level. In addition, the uncertainty of background estimation is minimized in this energy range since the predominant background component (internal background) is well understood, and the smearing of the source counts is minimal. While combining multiple archival observations may reduce statistical uncertainty in spectral shape and flux measurement, it can introduce even greater systematic uncertainties caused by instrumental (different telescope pointing) and physical (source variability) effects. We first grouped the archival observations taken at similar times (and hence similar position angles) together and performed spectral analysis within each cluster. We modeled the spectra with an absorbed power law where the hydrogen column density was fixed to $N_{H}=2.14\times 10^{22}$ cm-2 (Posselt & Pavlov, 2022). The abundance table from Wilms et al. (2000) was used. The model fits the data well (reduced $\chi^{2}$ $\sim$ 1). The best-fit power law indices were consistent between clusters within $1\sigma$ uncertainty. We then jointly analyzed all the archival observations with a cross-normalization term. The cross-normalization terms were measured to 1% precision. With respect to the focal plane module A (FPMA) spectrum of the oldest observation (observation ID 40001019002), the power-law normalization of the spectra was found to vary within $\pm$4range except for two outliers (observation ID 40021011002 and 40021015003, $-8$%). Therefore, we selected two representative archival observations (observation ID 40021002006 and 40021002008, total exposure 324 ks) to compare with the new observations, given the similar position angles between these observations. The two selected archival observations are referred to as the archival observations hereafter.
The best-fit power-law photon index for the archival observations is $3.42\pm 0.02$, and the 15-50 keV flux is $(4.73\pm 0.04)\times 10^{-11}$ erg s-1 cm-2. The new observations after 10 years show that the spectrum became marginally harder (power-law index $3.37\pm 0.02$), and the 15-50 keV flux ($(4.02\pm 0.04)\times 10^{-11}$ erg s-1 cm-2) decreased by $(15\pm 1)$(Figure 1). If the flux decreased simply by electrons’ synchrotron cooling, the spectrum should have become softer due to the inverse relation between synchrotron loss and electron energy. The observed marginal hardening of the X-ray spectrum, therefore, indicates that additional physical processes are operating in addition to synchrotron cooling.
@@ 3.1 SED model description
$\langle\dot{E}_{syn}\rangle=\frac{4}{3}\sigma_{T}\left(1-\frac{1}{\gamma}\right)\gamma^{2}\frac{B^{2}_{0}}{8\pi},$ (4)
@@ 4.1 Hard injection spectrum
Equivalently, the hard X-ray morphology locates only the current CR acceleration site (injected electrons), whereas the radio morphology also reveals the location of CRs accelerated in the past (preexisting electrons). The hard X-ray morphology of Cas A shows a high concentration of emission on the knots. On the other hand, the radio morphology is more isotropic with enhancement at the same knots and along the reverse shock (e.g. Domček et al. (2021)). The brightening in both wavelengths locates the knots as the current most active particle acceleration sites, while the globally isotropic radio emission traces the electrons that were accelerated at an earlier time, most likely at the forward shock, and diffused out.
@@ 4.2 Modified nonlinear diffusive shock acceleration
One can examine if the derived CR spectra are consistent with recent theoretical expectations. Nonlinear diffusive shock acceleration (NLDSA) theories address CR-driven modification to the standard DSA by introducing a region upstream of a shock with an enhanced density due to CR pressure (“precursor”, e.g. (Drury & Voelk, 1981; Drury, 1983; Jones & Ellison, 1991; Malkov & Drury, 2001)) that leads to an increased shock compression ratio ($R>4$). Recent hybrid (kinetic ion-fluid electron) simulations (Haggerty & Caprioli, 2020; Caprioli et al., 2020) discovered that the equivalent of a precursor is formed downstream of a shock (“postcursor”) whose contribution to the increased compression ratio dominates that of a precursor. On the other hand, the compression ratio experienced by CRs is decreased since magnetic fluctuations and CRs drift away from the shock in a postcursor. This postcursor effect on the “effective” compression ratio is parameterized by a factor $\alpha$, which depends on shock velocity, upstream density, and downstream magnetic field (see Methods). Within this framework (modified NLDSA or mNLDSA), the CR spectral index is given by $q=3R/(R-1-\alpha)-2$. Typical values of $R\sim 5$ and $\alpha\sim 0.6$ (Caprioli et al., 2020) predict $q\sim 2.4$ as we found for the preexisting electron population. At the reverse shock, our inferred spectrum of $q\sim 2.2$ is consistent with the theoretical prediction if the postcursor is not present (in which case $\alpha\rightarrow 0$ and $R\rightarrow 4$). This is plausible since, at the reverse shock, the postcursor may not form due to the presence of contact discontinuity in the downstream region.
In mNLDSA, the correction factor $\alpha$ for shock compression ratio can be deduced from observables as (Caprioli et al., 2020)
@@ 5 Summary and conclusion
Our multi-epoch hard X-ray observations enabled the isolation of a pure synchrotron radiation component of Cas A. Such radiation is associated with energetic, nonthermal CR electrons. By temporal modeling of this emission, we were able to establish the existence of two distinct populations of CR electrons. One population is more energetic and associated with a powerful, active accelerator, and the second population is less energetic and associated with an accelerator that was more active in the past. The X-ray morphology of Cas A allowed us to identify the sites of both accelerators. These observations and associated modeling provide the first self-consistent analysis of a young SNR that connects the CR spectrum and acceleration location to the broadband multiwavelength spectrum from radio to gamma-ray energies with interpretation through the most recent work on the mNLDSA theory. Our observational and theoretical approach can be applied to other young SNRs [^1] We are conducting hard X-ray variability studies of G1.9+0.3 and Kepler using multi-epoch NuSTAR observation data (papers in prep). A 500-ks NuSTAR observation of Tycho is scheduled in 2025 (PI: J.W.), providing a 10-year baseline with the NuSTAR observation in 2014. to elucidate the acceleration mechanism and environment of CRs below the “knee” of their spectrum at $\sim 3$ PeV.

### v2

@@ Abstract
Young supernova remnants (SNRs) are believed to be the origin of energetic cosmic rays (CRs) below the “knee” of their spectrum at $\sim 3$ petaelectronvolt (PeV, $10^{15}$ eV). Nevertheless, the precise location, duration, and operation of CR acceleration in young SNRs are open questions. Here, we report on multi-epoch X-ray observations of Cassiopeia A (Cas A), a 350-year-old SNR, in the 15-50 keV band that probes the most energetic CR electrons. The observed X-ray flux decrease ($15\pm 1\%$ over 10 years), contrary to the expected $>$90% decrease based on previous radio, X-ray, and gamma-ray observations, provides unambiguous evidence for CR electron acceleration operating in Cas A. A temporal model for the radio and X-ray data accounting for electron cooling and continuous injection finds that the freshly injected electron spectrum is significantly harder (exponential cutoff power law index $q=2.15$), and its cutoff energy is much higher ($E_{cut}=36$ TeV) than the relic electron spectrum ($q=2.44\pm 0.03$, $E_{cut}=4\pm 1$ TeV). Both electron spectra are naturally explained by the recently developed modified nonlinear diffusive shock acceleration (mNLDSA) mechanism. The CR protons producing the observed gamma rays are likely accelerated at the same location by the same mechanism as those for the injected electron. The Cas A observations and spectral modeling represent the first time radio, X-ray, gamma ray and CR spectra have been self-consistently tied to a specific acceleration mechanism -mNLDSA– in a young SNR.
@@ 2.1 Observations and data reduction
Table 1: List of archival and new observations
@@
40001019002 2012-08-18 84 0.9 291
40021001002 2012-08-27 76 2.0 170
40021001005 2012-10-07 33 1.5 183
40021002002 2012-11-23 338 0.3 271
40021002006 2013-03-02 249 3.9 136
40021002008 2013-03-05 249 4.0 189
40021003003 2013-05-28 151 3.5 198
40021011002 2013-10-30 7 1.8 236
40021012002 2013-11-27 335 1.1 206
40021015003 2013-12-23 312 1.6 137
Total exposure (archival data) 2,017
40801003002 2023-03-24 229 2.2 92
40801013002 2023-04-04 217 2.5 95
@@ 2.3 Spectral analysis
We analyzed the data in the 15-50 keV band. This choice of energy range ensures negligible contribution of thermal X-rays ($kT$ $<4$ keV (Hwang & Laming, 2012; Bleeker et al., 2001; Maeda et al., 2009b; Favata et al., 1997b; Allen et al., 1997b)) and source emission above the background level. In addition, the uncertainty of background estimation is minimized in this energy range since the predominant background component (internal background) is well understood, and the smearing of the source counts is minimal. While combining multiple archival observations may reduce statistical uncertainty in spectral shape and flux measurement, it can introduce even greater systematic uncertainties caused by instrumental (different telescope pointing) and physical (source variability) effects. To minimize these systematic uncertainties, we selected two representative archival observations (observation ID 40021002006 and 40021002008, total exposure 324 ks) to compare with the new observations, given their similar position angles to the new observations and the minimal time gap (3 days) between them. The two selected archival observations are referred to as the archival observations hereafter.
We extracted the source spectra for the entire remnant (redius $6^{\prime}$ circle). We modeled the spectra with an absorbed power law where the hydrogen column density was fixed to $N_{H}=2.14\times 10^{22}$ cm-2 (Posselt & Pavlov, 2022). The abundance table from Wilms et al. (2000) was used. The spectral parameters were linked among the observations within each epoch. A cross-normalization term was multiplied to each spectra with respect to the FPMA spectra of the earliest observations in each epoch. The average cross-normalizations are $1.019\pm 0.005$ for the archival data, and $1.011\pm 0.007$ for the new data. The model fits the data well (reduced $\chi^{2}$ $\sim$ 1). The best-fit power-law photon index for the archival observations is $3.42\pm 0.02$, and the 15-50 keV flux is $(4.73\pm 0.04)\times 10^{-11}$ erg s-1 cm-2. The new observations after 10 years show that the spectral shape remains unchanged within statistical uncertainties (power-law index $3.37\pm 0.02$), and the 15-50 keV flux ($(4.02\pm 0.04)\times 10^{-11}$ erg s-1 cm-2) decreased by $(15\pm 1)$(Figure 1). If the flux decreased simply by electrons’ synchrotron cooling, the spectrum should have become softer due to the inverse relation between synchrotron loss and electron energy. The null-detection of X-ray spectral softening, therefore, indicates that additional physical processes are operating in addition to synchrotron cooling.
For the readers who intend to combine the archival observations of Cas A into a single data point, we present the estimated systematic uncertainties on the spectral measurements for such a case [^1] While a detailed study of instrumental systematic uncertianties in spectral measurement was performed in Madsen et al. (2021) using the Crab nebula, its result is not directly applicable to our study of Cas A. The study in (Madsen et al., 2021) treated the Crab nebula as a point-like source and provided the systematic uncertainty due to a different off-axis angle in the range of 1-$7^{\prime}$ (1-$4^{\prime}$ for flux measurement). On the other hand, Cas A is an extended source with a highly inhomogeneous flux distribution; not only an off-axis angle but also a position angle becomes a source of systematic uncertainty in this case. Moreover, depending on the telescope pointing, part of Cas A ($\sim 6^{\prime}$ in diamaeter) may lie beyond the range of an off-axis angle covered by the study in Madsen et al. (2021) for flux measurement ($4^{\prime}$). Note also that the systematic uncertianties may include the intrinsic flux change of Cas A.. We analyzed all the archival observations with exposures greater than 100 ks (Table 1). We analyzed each observation separately, and linked the parameters between the FPMA and FPMB spectra of each observation using a cross-normalization term with respect to the FPMA. We report the standard deviations of the distributions of power-law index (0.6%), flux in 15-50 keV (4%), and cross-normalization (1.2%) as the systematic uncertainties in case the archival observations are combined into a single data point.
@@ 3.1 SED model description
$\langle\dot{E}_{syn}\rangle=\frac{4}{3}\sigma_{T}\left(1-\frac{1}{\gamma^{2}}\right)\gamma^{2}c\frac{B^{2}_{0}}{8\pi},$ (4)
@@ 4.1 Hard injection spectrum
Equivalently, the hard X-ray morphology locates only the current CR acceleration site (injected electrons), whereas the radio morphology also reveals the location of CRs accelerated in the past (preexisting electrons). The hard X-ray morphology of Cas A shows a high concentration of emission on the knots. On the other hand, the radio morphology is more isotropic with enhancement at the same knots and along the reverse shock (e.g. Domček et al. (2021)). The brightening in both wavelengths locates the knots as the current most active particle acceleration sites (e.g. Zirakashvili et al. (2014)), while the globally isotropic radio emission traces the electrons that were accelerated at an earlier time, most likely at the forward shock, and diffused out.
@@ 4.2 Modified nonlinear diffusive shock acceleration
One can examine if the derived CR spectra are consistent with recent theoretical expectations. Nonlinear diffusive shock acceleration (NLDSA) theories address CR-driven modification to the standard DSA by introducing a region upstream of a shock with an enhanced density due to CR pressure (“precursor”, e.g. (Drury & Voelk, 1981; Drury, 1983; Jones & Ellison, 1991; Malkov & Drury, 2001)) that leads to an increased shock compression ratio ($R>4$). Recent hybrid (kinetic ion-fluid electron) simulations (Haggerty & Caprioli, 2020; Caprioli et al., 2020) discovered that the equivalent of a precursor is formed downstream of a shock (“postcursor”) whose contribution to the increased compression ratio dominates that of a precursor. On the other hand, the compression ratio experienced by CRs is decreased since magnetic fluctuations and CRs drift away from the shock in a postcursor. This postcursor effect on the “effective” compression ratio is parameterized by a factor $\alpha$, which depends on shock velocity, upstream density, and downstream magnetic field. In mNLDSA, the correction factor $\alpha$ for shock compression ratio can be deduced from observables as (Caprioli et al., 2020)
@@
Within this framework (modified NLDSA or mNLDSA), the CR spectral index is given by $q=3R/(R-1-\alpha)-2$. Typical values of $R\sim 5$ and $\alpha\sim 0.6$ (Caprioli et al., 2020) predict $q\sim 2.4$ as we found for the preexisting electron population. At the reverse shock, our inferred spectrum of $q\sim 2.2$ is consistent with the theoretical prediction if the postcursor is not present (in which case $\alpha\rightarrow 0$ and $R\rightarrow 4$). This is plausible since, at the reverse shock, the postcursor may not form due to the presence of contact discontinuity in the downstream region.
@@ 5 Summary and conclusion
Our multi-epoch hard X-ray observations enabled the isolation of a pure synchrotron radiation component of Cas A. Such radiation is associated with energetic, nonthermal CR electrons. By temporal modeling of this emission, we were able to establish the existence of two distinct populations of CR electrons. One population is more energetic and associated with a powerful, active accelerator, and the second population is less energetic and associated with an accelerator that was more active in the past. The X-ray morphology of Cas A allowed us to identify the sites of both accelerators. These observations and associated modeling provide the first self-consistent analysis of a young SNR that connects the CR spectrum and acceleration location to the broadband multiwavelength spectrum from radio to gamma-ray energies with interpretation through the most recent work on the mNLDSA theory. Our observational and theoretical approach can be applied to other young SNRs to elucidate the acceleration mechanism and environment of CRs below the “knee” of their spectrum at $\sim 3$ PeV.